OpenStudy (anonymous):

check my answer? medal! c:

OpenStudy (anonymous):

Solve each system by substitution 11. { 14x – 35 = 7y -25 – 6x = 5y 14x – 35 = 7y 7(2x – 5) = 7y y= 2x – 5 -25 – 6x = 5y -6x – 5y – 25 = 0 y=(-6x – 25) (-4,-7.5) Solve each system by elimination. 13. { 2x + 3y = 4 4x + 6y = 9 2x = -3y + 4 x = (-3y+4)/2 4[(-3y+4)/2] - 9y = 9 2(-3y+4) - 9y = 9 -6y + 10 - 9y = 9 -15y = -1 y = 1/15 x = (-3y+4)/2 x = (-3(1/15)+4)/2 x = [(-1/5)+4/2 = (24/4)/2 = 12/5 Final Answer: y = -1/3 when x = 12/5

OpenStudy (anonymous):

oops can you see the whole equation for #11?

jimthompson5910 (jim_thompson5910):

I'm getting different answers for #11 and #13

jimthompson5910 (jim_thompson5910):

when you get to -6x – 5y – 25 = 0, plug in y = 2x – 5 and solve for x -6x – 5y – 25 = 0 -6x – 5(2x-5) – 25 = 0 x = ???

OpenStudy (anonymous):

oh dang :/ I thought I finally knew how to solve these problems.. do you mind helping me? showing how you got your answers im so tired ive been working on these problems since 2 pm ugh

jimthompson5910 (jim_thompson5910):

for #11, you got -6x – 5y – 25 = 0 correct

jimthompson5910 (jim_thompson5910):

now you plug in y = 2x - 5 and solve for x

jimthompson5910 (jim_thompson5910):

so you have to solve -6x – 5(2x-5) – 25 = 0

OpenStudy (anonymous):

for #11 I got { 14x–35=7y -25–6x=5y 14x–35=7y 7(2x–5)=7y y=2x–5 -25–6x=5y -6x–5y–25=0 y=(-6x–25) (-4,-7.5)

OpenStudy (anonymous):

all I am seeing are boxes?

jimthompson5910 (jim_thompson5910):

refresh the page

jimthompson5910 (jim_thompson5910):

the boxes will go away

OpenStudy (anonymous):

okay got it :)

OpenStudy (anonymous):

they went away s:

jimthompson5910 (jim_thompson5910):

are you able to solve -6x - 5(2x-5) - 25 = 0

OpenStudy (anonymous):

x=0?

jimthompson5910 (jim_thompson5910):

then you will use that value of x to find y

jimthompson5910 (jim_thompson5910):

x = 0 y = 2x - 5 y = ???

OpenStudy (anonymous):

y=2x-5=0

jimthompson5910 (jim_thompson5910):

if x = 0, then y is ... y = 2x - 5 y = 2(0) - 5 ... replace every x with 0 y = ???

OpenStudy (anonymous):

y=-5?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so the answer to #11 is (0, -5)

jimthompson5910 (jim_thompson5910):

since you found that x = 0 and y = -5

OpenStudy (anonymous):

thank you soooo much!! :) do you mind helping me with #13 also?

jimthompson5910 (jim_thompson5910):

notice your second equation in #13 is 4x + 6y = 9

jimthompson5910 (jim_thompson5910):

yet you somehow have 4[(-3y+4)/2] - 9y = 9 after you plugged in that x value. Where did the -9y come from?

jimthompson5910 (jim_thompson5910):

it should be 4[(-3y+4)/2] + 6y = 9

OpenStudy (anonymous):

I honestly don't know my friend was trying to help me :c

jimthompson5910 (jim_thompson5910):

I see

OpenStudy (anonymous):

sorry I suck at math /: any other subject im good at lol

jimthompson5910 (jim_thompson5910):

thats ok

jimthompson5910 (jim_thompson5910):

are you able to compute 4[(-3y+4)/2]

jimthompson5910 (jim_thompson5910):

or simplify that

OpenStudy (anonymous):

yes hold on 2(4-3y)?

jimthompson5910 (jim_thompson5910):

keep going

OpenStudy (anonymous):

simplify that answer also?

jimthompson5910 (jim_thompson5910):

distribute

OpenStudy (anonymous):

8-6y?

jimthompson5910 (jim_thompson5910):

4[(-3y+4)/2] + 6y = 9 turns into 8-6y + 6y = 9

OpenStudy (anonymous):

thank you c: do I keep going? or is that the answer?

jimthompson5910 (jim_thompson5910):

keep going and try to solve for y

OpenStudy (anonymous):

y= 0?

jimthompson5910 (jim_thompson5910):

what is -6y + 6y

OpenStudy (anonymous):

I keep getting 0 on my calculator... 36y?

jimthompson5910 (jim_thompson5910):

maybe think of it as 6y - 6y

jimthompson5910 (jim_thompson5910):

but yeah you get 0y or just 0

jimthompson5910 (jim_thompson5910):

8-6y + 6y = 9 turns into 8+0 = 9 and that becomes 8 = 9

jimthompson5910 (jim_thompson5910):

but I'm sure you see the problem

OpenStudy (anonymous):

thank you again I really appreciate you helping me! c:

jimthompson5910 (jim_thompson5910):

so what's the answer to #13?

OpenStudy (anonymous):

(9,0)?

jimthompson5910 (jim_thompson5910):

is 8 = 9 true?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

since it's always false, regardless of what x and y are, this means that the overall system has no solutions at all. There's no way to make that equation true.

jimthompson5910 (jim_thompson5910):

visually, you'll have 2 parallel lines that never intersect you need a point of intersection to have a solution

OpenStudy (anonymous):

ohh duh >_< sorry wasn't thinking correctly